### v1

Abstract
There is growing evidence from gravitational-wave observations that some merging black holes are created from previous mergers. Using the prediction that these hierarchically-merged black holes have dimensionless spin magnitudes of $\chi\approx 0.69$, we identify a subpopulation in the gravitational-wave data consistent with a hierarchical-merger origin in dense star clusters. This subpopulation’s primary mass distribution peaks at $16.7^{+3.1}_{-4.6}\,\mathrm{M}_{\odot}$, which is approximately twice as large as its secondary mass distribution’s mode ($8.0^{+29.7}_{-2.3}\,\mathrm{M}_{\odot}$), and its spin tilt distribution is consistent with isotropy. Our inferred secondary mass distributions imply that isolated binary evolution may still be needed to explain the entirety of the $9\,\mathrm{M}_{\odot}$ peak. Surprisingly, we find that the rate of hierarchical mergers may evolve more steeply with redshift than the rest of the population ($98.5\%$ credibility): the fraction of all binary black holes that are hierarchically formed at $z=0.1$ is $0.05^{+0.05}_{-0.04}$, compared to $0.17^{+0.13}_{-0.12}$ at $z=1$. This provides an explanation for the previously-discovered broadening of the effective spin distribution with redshift. Our results have implications for star cluster formation histories, as they suggest the potential existence of a high-redshift population of massive, compact clusters.
[…]
software: gwpopulation (Talbot et al., 2025b), gwpopulation_pipe (Talbot, 2021), Dynesty (Speagle, 2020), Jupyter (Perez and Granger, 2007; Kluyver et al., 2016), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (McKinney, 2010; pandas development team, 2025), python (Van Rossum and Drake, 2009), scipy (Virtanen et al., 2020; Gommers et al., 2025), Bilby (Ashton et al., 2019; Romero-Shaw et al., 2020; Talbot et al., 2025a), corner.py (Foreman-Mackey, 2016; Foreman-Mackey et al., 2024), Cython (Behnel et al., 2011), JAX (Bradbury et al., 2018), and tqdm (da Costa-Luis et al., 2024).
I Introduction
The stellar-mass binary black holes detectable by LIGO (LIGO Scientific Collaboration et al., 2015; Capote et al., 2025; Soni et al., 2025), Virgo (Acernese et al., 2015) and KAGRA (Akutsu et al., 2021) (LVK) are likely created from the collapse of massive stars. These stars may be born as a binary system in the galactic field that then evolves into a binary black hole system, or they may be born in a dense stellar environment such as a star cluster, in which the black hole stellar remnants dynamically assemble into tightly-bound binaries, and typically encounter other objects in the cluster before merger (see reviews by Mapelli, 2020; Mandel and Farmer, 2022, and references therein). While it is not yet known which of these potential formation scenarios, if any, contributes to the majority of binary black hole systems, one clear signature of the dynamical assembly hypothesis is the existence of hierarchical mergers (e.g. Gerosa and Fishbach, 2021). Hierarchical mergers include one or more “second-generation” remnants of previous, “first-generation” black hole mergers, and are informative for a number of reasons.
Firstly, they are rarely produced in clusters with low stellar mass, as the recoil kicks that “second-generation” black holes receive from anisotropic emission of gravitational waves are typically larger than the clusters’ escape velocity. Thus, “second-generation” black holes are often ejected from star clusters and only a small fraction remain to undergo another merger  (Fitchett, 1983; Portegies Zwart and McMillan, 2000; Favata et al., 2004; González et al., 2007; Lousto et al., 2010; Gerosa et al., 2018; Mahapatra et al., 2021). The mass and spin of the “second-generation” object in a hierarchically-merged system therefore provide a lower limit on the escape velocity of its host and insight into the system’s environment.
Figure 1: The inferred merger rate evolution with redshift of our subpopulation of $\chi_{1}\approx 0.7$ binary black holes ($p_{\rm sub}(z|\kappa_{\rm sub})$, green in left panel) compared to that of the rest of the population ($p_{\rm main}(z|\kappa_{\rm main})$, violet in left panel). Individual draws from the main and subpopulation are shown as thin lines, shaded bands show the $90\%$ credible interval, and dark lines show the mean. As shown by the inferred power law indices for the main and subpopulation (right panel), the spinning subpopulation has a steeper redshift distribution than the main population, causing a higher contribution to the overall rate (orange lines in left panel) at high redshift than at low redshift. This is contrary to typical expectations of hierarchical mergers in star clusters, which necessitate slightly longer delay times for hierarchical mergers than the first-generation mergers that precede them. The existence of a rapidly-spinning subpopulation with a steep redshift evolution provides a natural explanation of the previously-identified correlation between effective spin and redshift in the overall binary black hole population.
Secondly, if assumptions about the typical star clusters that host dynamically-formed binary black holes are made, the number of mergers that include a “second-generation” black hole strongly informs the number of dynamically-formed “first-generation”+“first-generation” mergers in the population (Kimball et al., 2020), thereby constraining the total rate of dynamically assembled systems.
Finally, hierarchical mergers have clear population predictions that allow them to be straightforwardly distinguished from the larger binary black hole population (Fishbach et al., 2017; Gerosa and Berti, 2017). An ensemble of black holes that have resulted from previous mergers will have a spin magnitude distribution peaked near $\chi=0.69$ (Berti and Volonteri, 2008; Hofmann et al., 2016; Rodriguez et al., 2019; Borchers et al., 2025), regardless of how many previous mergers occurred (Fishbach et al., 2017). Additionally, “second-generation”+“first-generation” mergers will have a mass ratio distribution peaked near $\approx 1/2$ (Gerosa and Berti, 2017; Rodriguez et al., 2019; Kimball et al., 2020), and binary black holes formed in gas-free star clusters should have an isotropic spin tilt distribution (but cf. Kıroğlu et al. 2025). However, mass ratio and spin orientation distributions differ for 2G+1G mergers formed in the disks of active galactic nuclei (Tagawa et al., 2020; McKernan et al., 2020; Vaccaro et al., 2024; Vaccaro, 2025; Li et al., 2025b).
The simplicity and robustness of these predictions contrasts with most other formation pathways whose population predictions suffer from large theoretical uncertainties. Hierarchical mergers may therefore be the first subpopulation to be confidently identified, even if they make up a small fraction of the total population. Indeed, evidence for a hierarchically-merged subpopulation in the gravitational wave data is growing. In addition to hints that hierarchical mergers are filling the pair-instability mass gap in the binary black hole primary mass distribution (Pierra et al., 2024; Antonini et al., 2025; Tong et al., 2025b; Afroz and Mukherjee, 2025), the two gravitational wave detections recently released by the LVK with primary spin magnitudes of $\approx 0.7$ (Abac et al., 2025a) have pointed to hierarchical mergers across the binary black hole mass spectrum (Tong et al., 2025a; Berti et al., 2025), a possibility first pointed out by Tiwari and Fairhurst (2021).
In this work, we isolate the hierarchical mergers in the latest gravitational wave data using a phenomenological model that takes advantage of the most straightforward and robust population prediction for hierarchical mergers: their primary spin magnitude distribution. We then investigate the population properties of this hierarchically-merged subpopulation and provide insight into their cluster environments and the population of “first-generation” black holes in star clusters (via the 1G secondaries of 2G+1G mergers). The redshift distribution of our identified hierarchically-merged subpopulation is likely steeper than that of the rest of the population (Fig. 1). This is in contrast to naïve expectations from cluster dynamics, in which hierarchical mergers occur at similar (or slightly later) times in a cluster’s history since they necessarily occur after the first “first-generation” mergers. This finding therefore either implies that the subpopulation of $\chi_{1}\approx 0.69$ objects is not formed hierarchically, or that the hierarchical mergers are coming from a distinct population of more distant star clusters than produced our observed first-generation mergers. Regardless, the fact that a highly-spinning subpopulation has a steeper redshift evolution than the rest of the binary black holes offers a potential explanation to the previously-discovered broadening of the effective spin distribution with redshift. In a companion paper (Vijaykumar et al., in prep.), we show that this subpopulation can also explain the effective spin distribution’s narrowing with mass ratio (Callister et al., 2021; The LIGO Scientific Collaboration et al., 2025).
The low-mass end of our inferred 1G mass distribution differs between the 1G+1G and 2G+1G populations by more than would be expected from the simplest mass segregation models. This suggests that other formation channels, such as isolated binary evolution, contribute to the low-mass 1G+1G population.
This paper is organized as follows. Section II describes the model we use to isolate a potential subpopulation of “second-generation”+“first-generation” binary black holes, and we present the evidence for such a subpopulation in Section III. Section IV describes our findings related to the redshift evolution of this subpopulation, using it to provide measurements of the relative fraction of cluster-origin binary black holes in the overall population. Our inferred mass distributions are shown in Section V. Section VI interprets our results in the context of previously-identified correlations in the population, including correlations between primary mass and spin and correlations between spin and redshift, and Section VII discusses their implications for cluster properties and the contribution of the isolated binary channel to the overall BBH merger rate. The appendices provide additional results, such as the distributions of spin magnitudes, spin orientations, and mass ratios for the subpopulation and main population.
II Phenomenological population model
We construct a phenomenological mixture model to search for a subpopulation, $S$, of 2G+1G hierarchical mergers in the GW data and disentangle it from the main population, $M$. The subpopulation is defined by its primary spin magnitude distribution, which is peaked at $\chi_{1}=0.69$. We emphasize that although this is a strong prior choice designed to isolate a population of 2G+1G mergers, we do not enforce the existence of this subpopulation. The subpopulation’s other dimensions (i.e. its mass, spin orientation, and redshift distributions) are fit separately from the main population. The mixture model therefore has the form
$p(m_{1},\cos{\theta_{1}},\chi_{1},m_{2},\cos{\theta_{2}},\chi_{2},z|\Lambda)$ $=\xi p_{S}(m_{1},m_{2},\cos{\theta_{1}},\cos{\theta_{2}},z|\lambda_{S})\mathcal{N}(\chi_{1}|\mu^{\chi_{1}}_{S},\sigma^{\chi_{1}}_{S})\mathcal{N}(\chi_{2}|\mu^{\chi_{2}}_{S},\sigma^{\chi_{2}}_{S})$ (1)
$+(1-\xi)p_{M}(m_{1},m_{2},\cos{\theta_{1}},\cos{\theta_{2}},z|\lambda_{M})\mathcal{N}(\chi_{1}|\mu^{\chi},\sigma^{\chi})\mathcal{N}(\chi_{2}|\mu^{\chi},\sigma^{\chi}),$
where $\xi$ is the fraction of systems in $S$. Here, $m_{i}$, $\theta_{i}$, and $\chi_{i}$ are the component masses, spin tilt angles, and spin magnitudes, respectively. $\Lambda$ is the set of all hyperparameters, $\lambda_{j}$ are the sets of hyperparameters that govern the distribution of masses and spin tilts for both subpopulations, and $j=\{S,M\}$. Gaussian distributions with mean $\mu$, standard deviation $\sigma$ and truncated in the range $[0,1]$ are represented by $\mathcal{N}(\cdot|\mu,\sigma)$. Both populations use a pairing function formalism for the mass distribution, allowing us to fit for separate component mass distributions as well as their mass-ratio dependent pairing (Fishbach and Holz, 2020). We describe the forms of the main and subpopulation in more detail in Appendix A.
Using a hierarchical Bayesian analysis (Loredo, 2004; Thrane and Talbot, 2019; Vitale et al., 2022b), we fit this model to all BBHs in GWTC-4, which includes BBHs in all previous catalogs. We additionally include GW241110 and GW241011 (Abac et al., 2025a), two events that have been recently published by the LVK because their spins and mass ratios are consistent with a hierarchical merger origin. However, excluding these events does not noticeably change our findings. We omit GW231123  (Abac et al., 2025b) as its large secondary mass (above $100\,M_{\odot}$) appears to be an outlier with respect to our inferred secondary mass distributions. Including this event does not affect our redshift or spin inferences. Appendix B further discusses the effects of GW231123 on our inferred mass distributions.
III Identification of a $\chi_{1}\approx 0.7$ subpopulation
We find strong evidence for a subpopulation of $\chi_{1}=0.69$ systems. The fraction of systems in this subpopulation, $\xi$, is larger than $0.02$ at the $99.7\%$ level, representing a Bayes factor of $20.3$. This subpopulation has spin orientation, mass ratio, and component mass distributions characteristic of hierarchical mergers in star clusters (see Appendices C.1, D, and Section V, respectively).
Marginalized over the range $z=[0,2]$, our inferred value of $\xi=0.23^{+0.18}_{-0.17}$ is roughly consistent with previous studies that also identify a 2G+1G-like subpopulation that makes up roughly $12-20\%$ of the population [^1] Antonini et al. (2025) estimate that $20\%$ of all mergers are 2G+1G, and Tong et al. (2025a) estimate the rate of 2G+1G mergers to be $2.2^{+1.9}_{-1.2}\,\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, which corresponds to $\approx 12\%$ of the total BBH merger rate inferred by The LIGO Scientific Collaboration et al. (2025), albeit with a different population model. Kimball et al. (2021) find $\mathcal{R}_{\rm 2G+1G}/\mathcal{R}_{\rm 1G+1G}<0.12$ when analyzing GWTC-2, and our posterior on this value peaks at $0.23$, putting these two studies performed on different datasets in mild tension. (e.g. Kimball et al., 2021; Wang et al., 2022; Mould et al., 2022; Pierra et al., 2024; Antonini et al., 2025; Tong et al., 2025b, a; Berti et al., 2025; Vijaykumar et al., in prep.). The fact that similar fractions of 2G+1G systems are inferred – regardless of what compact binary parameter (i.e. effective spin, mass ratio, or primary spin magnitudes) is used to identify them – strengthens the evidence that these independently-discovered subpopulations are indeed the same as one another and represent 2G+1G mergers.
The inferred fraction of 2G+1G mergers is high when marginalized over the full redshift range in which we define our model. Taken at face value, this may lead to the conclusion that all observed compact binaries can be sourced from clusters, as $\mathcal{O}(10\%)$ of all cluster mergers should be hierarchical (Rodriguez et al., 2019). However, we find that the fraction of hierarchical mergers depends strongly on redshift, with low fractions in the local Universe. Thus, the local ($z\lesssim 0.3$) merger rate is difficult to explain with cluster mergers alone, and further study into the redshift-dependent mass function of star clusters is necessary to compare with the observed gravitational wave population.
IV The hierarchical merger rate evolves faster with redshift than the rest of the population
The merger rate of the $\chi_{1}\approx 0.7$ subpopulation may evolve more steeply than that of the main population. Figure 1 shows the merger rate evolution of both subpopulations as a function of redshift, as well as the posteriors on the power law index of each population’s redshift distributions, $\kappa_{\rm main}$ and $\kappa_{\rm sub}$. The power law index of the total population, as inferred by The LIGO Scientific Collaboration et al. (2025), is $3.2^{+0.94}_{-1.00}$, which lies between the values we infer for our main and subpopulations. The posterior on $\kappa_{\rm main}$ is less than that on $\kappa_{\rm sub}$ at $98.5\%$ credibility. There are therefore hints, but not yet definitive evidence, that hierarchical mergers represent a higher fraction of the binary black hole population at $z=1$ compared to $z=0$. If it is a real feature of the population, this result can explain the previously-observed broadening of the effective spin distribution with redshift (Biscoveanu et al., 2022), as we discuss in Section VI.
Hierarchical mergers necessarily proceed the first 1G mergers in a cluster, and therefore occur at similar or slightly later times in a clusters’ history than 1G mergers do. This expectation appears to be in contradiction with the results shown in Figure 1, implying either that our subpopulation is not a set of 2G+1G mergers (in contrast to our spin and mass ratio results), or that a more complex picture of star cluster contributions to the BBH merger rate is necessary to explain the data. Assuming a cluster mass function that scales inversely with the square of the cluster mass, the metallicity-specific star formation history from Madau and Fragos (2017), and the catalog of Cluster Monte Carlo simulations from Kremer et al. (2020), we infer the fraction of all BBHs that originated in star clusters to be $0.32^{+0.35}_{-0.23}$ at $z=0.1$ and $1.23^{+0.93}_{-0.84}$ at $z=1$. This directly follows from our measured fraction of 2G+1G mergers in the total population at different redshifts, which is $0.05^{+0.05}_{-0.04}$ at $z=0.1$ and $0.17^{+0.13}_{-0.12}$ at $z=1$. Thus, if our assumptions about the redshift-dependent mass function of clusters are accurate, contributions from other formation scenarios may be necessary to explain the merger rate at low redshift, but clusters can explain the full merger rate at high redshift. Additionally, the fact that the fraction of cluster-origin BBHs may be larger than unity at $z=1$ means that denser clusters at high redshift may be required to explain our inferred fraction of 2G+1G mergers. Theoretical uncertainties on cluster properties as a function of redshift remain uncertain, and future work will aim to constrain these with the subpopulation’s inferred redshift distribution (e.g. Fishbach and Fragione, 2023). We further discuss the potential astrophysical implications of our inferred redshift distributions in Section VII.
V Mass distributions
Figure 2: Inferred primary (top panel) and secondary (bottom panel) mass distributions for the hierarchically-merged subpopulation (green) and main population (violet), after a pairing function is removed. The primary mass distribution of the subpopulation peaks at roughly twice the value of its secondary mass distribution’s mode. Additionally, only the subpopulation’s primary mass distribution contributes significantly to the merger rate at high masses. These two facts imply that the $\chi_{1}\approx 0.7$ subpopulation is primarily made up of “second-generation”+“first-generation” mergers, whereas “first-generation”+“first-generation” mergers (originating either in star clusters or in the galactic field) constitute the main population. It is therefore possible to interpret the subpopulation’s primary and secondary mass distributions as the distributions of “second-generation” and “first-generation” black holes in clusters, respectively.
Our inferred component mass distributions for each subpopulation are shown in Fig. 2. As in Farah et al. (2024), we display the underlying component mass distributions (i.e. after the pairing function has been factored out), but we provide marginalized component mass distributions in Appendix F. The hierarchical subpopulation’s primary mass distribution (representing 2G black holes) has a mode at $16.7^{+3.1}_{-4.6}\,\mathrm{M}_{\odot}$, which may be approximately twice as large as the mode of its secondary mass distribution ($8.0^{+29.7}_{-2.3}\,\mathrm{M}_{\odot}$). This is consistent with a 2G+1G interpretation of the subpopulation.
Additionally, the subpopulation contributes the majority of the merger rate at high primary masses ($\gtrsim 60\,\mathrm{M}_{\odot}$), as shown by the agreement between the orange (total) and green (subpopulation) bands in that region. The subpopulation’s secondary mass distribution does not extend to high masses, whereas the main population’s primary and secondary mass distributions appear to be similar. Even though their maximum masses are allowed to differ, the primary and secondary mass distributions of the main (1G+1G) population both exhibit negligible merger rates above $\approx 46\,\mathrm{M}_{\odot}$ (see Fig. 12 in Appendix F). Taken together, these results further support the hypothesis proposed in Tong et al. (2025b) that the pair-instability mass gap is polluted primarily by 2G+1G mergers, and not by 2G+2G mergers. However, it is possible that “second-generation”+“second-generation” mergers are present but not yet distinguishable from the rest of the population, and we cannot rule out their presence.
The subpopulation’s primary mass distribution need not have a secondary mode at higher masses, whereas the main population’s primary mass distribution is clearly bimodal, with a secondary mode at $\approx 35\,\mathrm{M}_{\odot}$. This creates a “shelf” rather than a peak in the total primary mass distribution, which has been observed by The LIGO Scientific Collaboration et al. (2025).
Our inferred mass distributions are qualitatively consistent with those found by previous studies that have searched for correlations between mass and spin (Biscoveanu et al., 2022; Pierra et al., 2024; Tiwari, 2025; Tong et al., 2025a; Banagiri et al., 2025; Berti et al., 2025).
[…]
This implies that the $9\,\mathrm{M}_{\odot}$ peak cannot be entirely explained by the dynamical formation hypothesis, unless the “first-generation” black hole population in star clusters differs significantly between those that participate in “first-generation”+“first-generation” mergers versus “second-generation”+“first-generation” mergers. This difference goes beyond what would already be produced by a mass ratio-dependent pairing function, which naturally favors more equal-mass pairings (and therefore selects more massive “first-generation” black holes to participate in “second-generation”+“first-generation” mergers), as Figs. 2 and 3 use the component mass distributions with the pairing function factored out. Such a pairing function should account for mass segregation [^2] It is possible that our mass ratio-dependent pairing function is an insufficient descriptor of BBH pairing in clusters. For example, if BBH pairing depends on non-mass-ratio paramters, or if the pairing mechanism differs between 2G+1G versus 1G+1G mergers, as might be expected if BH ejections and other cluster dynamics cause the pairing function to change over the clusters’ histories.. Therefore, isolated evolution may be needed to explain the low-mass end of the observed population. This is consistent with the spin orientation distribution of the main “first-generation”+“first-generation” population, which has an aligned-spin component (see Appendix C.1), and may therefore require a contribution from active galactic nuclei or isolated binary evolution.
Figure 3: Ratio of the subpopulation’s “first-generation” mass distribution, representing “first-generation” BHs that merge in clusters, to the main population’s mass distribution, which consists of “first-generation” BHs in clusters as well as all other binary formation scenarios. Horizontal lines on this plot would indicate that all binary black holes originate in the same population of star clusters that produce 2G+1G mergers. However, a noticeable dip is present at $\approx 9\,\mathrm{M}_{\odot}$, implying either that isolated binary evolution is required to fully explain the feature there, or that the clusters that produce our observed 1G+1G mergers differ from those that produce the 2G+1G population.
VI Hierarchical Mergers Can Source Correlations Between Effective Spin, Redshift, Mass, and Mass Ratio
Figure 4: Population-level correlations caused by hierarchical mergers. Diagonal panels show the inferred one-dimensional mass, effective spin, and redshift distributions of the main population (violet), 2G+1G subpopulation (green) and combined, or total population (orange). These are marginalized over all other dimensions. Off-diagonal panels show filled contours of the total population in orange, with locations of the main and subpopulation contributions indicated with unfilled dashed and solid contours, respectively. All two-dimensional contours show hyperposterior-averaged distributions, while the one-dimensional mass and redshift distributions show hyperposterior uncertainty as shaded bands. Correlations can be observed between all parameters shown: the high-mass tail of the population increases with redshift, the effective spin distribution broadens with redshift, and the effective spin distribution modulates with mass.
Fig. 4 summarizes the results of our analysis in the context of population-level correlations. Note that the effective spin ($\chi_{\rm eff}$) distribution cannot be analytically calculated from our component spin model, so it is reconstructed from samples and is therefore shown as a histogram, marginalized over hyperparameter uncertainty. All pairs of parameters shown in Fig. 4 exhibit correlations in the total population (orange contours), even though the correlations are absent in the individual subpopulations. In particular, the hierarchical subpopulation’s high primary spin magnitudes and isotropic spin tilts (see Appendix C) cause it to have a broader $\chi_{\rm eff}$ distribution than our main population (center panel). Its low contribution to the merger rate at low redshift but steep increase in contribution at high redshift therefore causes the overall-$\chi_{\rm eff}$ distribution to broaden with redshift (bottom center panel), offering an explanation to the correlation first discovered by Biscoveanu et al. (2022).
In a companion paper (Vijaykumar et al., in prep.), we show that a subpopulation of hierarchical mergers also offers a natural explanation of the correlation first discovered in Callister et al. (2021), and then found by The LIGO Scientific Collaboration et al. (2025) to be a narrowing of the effective spin distribution with mass ratio in GWTC-4 (but see Li et al., 2025a, for an explanation from active galactic nuclei). All previously-identified correlations in the GW data can therefore be attributed to a single origin.
Thus, binary black holes may be suffering from inverse Simpson’s paradox (Simpson, 1951), whereby the trends observed in the aggregate data (i.e. correlations between effective spin, redshift, and mass ratio in the full binary black hole population) disappear or reverse when the data is considered in separate groups (in this case, 2G+1G versus 1G+1G BBHs).
An important caveat to our interpretation is that if the broadening of the $\chi_{\rm eff}$ distribution with redshift were not caused by hierarchical mergers, our model might spuriously identify it as a steeper redshift dependence of a highly-spinning subpopulation relative to the main population. One proposed explanation is the correlation between tidal spin-up and delay-time in binary star systems, as both depend on orbital separation (Bavera et al., 2022), but we disfavor this explanation as our subpopulation appears to have an isotropic spin tilt distribution. Regardless of the interpretation, however, our findings reinforce the fact that spin is correlated with redshift on a population level.
VII Discussion
Using the prediction that second-generation BHs should have dimensionless spin magnitudes of $\chi\approx 0.69$, we identify a subpopulation of hierarchical mergers in the fourth gravitational-wave transient catalog. The merger rate and primary mass distribution of this subpopulation is consistent with those found previously with alternative methods (Tiwari and Fairhurst, 2021; Mould et al., 2022; Wang et al., 2022; Pierra et al., 2024; Antonini et al., 2025; Tiwari, 2025; Tong et al., 2025a; Vijaykumar et al., in prep.). Having isolated this hierarchically-formed subpopulation, we study its component mass, spin orientation, and redshift distributions.
We find that the rate of hierarchical mergers likely evolves more steeply with redshift than the rest of the population. This would imply that 2G+1G mergers occur at earlier times than many of the the 1G+1G mergers observable in current gravitational wave data. This is potentially surprising, as hierarchical mergers cannot occur before “second-generation” black holes are made through “first-generation”+“first-generation” mergers. Ye and Fishbach (2024) shows that if the cluster mass function does not evolve with redshift, no significant difference is expected between the redshift evolution of 1G+1G versus hierarchical mergers (except at very early times before any “second-generation” black holes have formed). Our results that hierarchical mergers tend to occur at higher redshifts than “first-generation”+“first-generation” mergers, and that the fraction of hierarchical mergers may be as high as $0.17^{+0.13}_{-0.12}$ at $z=1$, may therefore imply that the cluster mass function does in fact evolve with redshift, and that hierarchical mergers originate in star clusters that are denser, more massive, and at higher redshift than the lower-redshift population that hosts the rest of the observed mergers (e.g. Mai et al., 2025). This interpretation is consistent with recent near-infrared observations, which suggest that high-redshift clusters are formed with higher densities, masses, and occurence rates than the low-redshift clusters against which many cluster simulation codes are calibrated (Vanzella et al., 2022; Mowla et al., 2022; Adamo et al., 2024; Mowla et al., 2024; Mayer et al., 2025).
Another potential explanation for the steeper redshift evolution of a hierarchical subpopulation is that isolated binary stars source the observed low-redshift mergers. These systems may have longer delay times between star formation and merger, or may have been formed later in the Universe’s history than stars born in clusters. The latter is a natural consequence of the fact that the star formation rate in clusters peaks at higher redshifts than the total star formation rate, even when accounting for the low metallicities that source massive black holes. This hypothesis may be favored if the $9\,\mathrm{M}_{\odot}$ peak has a shallower redshift evolution than the rest of the mass distribution, as Fig. 3 suggests that non-star cluster channels contribute to this feature.
The steeper redshift evolution of the 2G+1G population when compared to the 1G+1G population offers an explanation for the effective spin distribution’s broadening with redshift (Biscoveanu et al., 2022). If all observed correlations in the gravitational wave data can be attributed to a subpopulation of hierarchical mergers (see e.g. Vijaykumar et al., in prep.), a consistent picture of the binary black hole population is beginning to emerge. This allows us to begin interpreting the rest of the binary black hole population, using the 2G+1G subpopulation as an anchor. Furthermore, the inferred population properties of hierarchical mergers allows us to understand the properties of star clusters from which they are sourced.
The authors gratefully acknowledge helpful conversations with Hui Tong, Daniel Holz, Utkarsh Mali, Mike Zevin, Sharan Banajiri, and Thomas Dent. We extend special thanks to Bart Ripperda for granting us access to the Bee workstation at CITA, where the computations for this work were performed. AV acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) (funding reference number 568580). MF acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) under grant RGPIN-2023-05511, the Alfred P. Sloan Foundation, and the Ontario Early Researcher Award. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). This research has made use of the Astrophysics Data System, funded by NASA under Cooperative Agreement 80NSSC21M00561.
Appendix A Population model details
The functional forms for the main population and subpopulation are similar, though all parameters of these models differ between the two. Following The LIGO Scientific Collaboration et al. (2025), we model the spin tilt distribution for both populations as a mixture between an isotropically-distributed component and an aligned-spin component. Within each population, the component spin tilts – $\cos{\theta_{1}}$ and $\cos{\theta_{2}}$, are identically and independently distributed. The main population is required to have independent but identically distributed spin magnitudes ($\mu^{\chi_{1}}_{M}=\mu^{\chi_{2}}_{M}=\mu^{\chi}_{M}$ and $\sigma^{\chi_{1}}_{M}=\sigma^{\chi_{2}}_{M}=\sigma^{\chi}_{M}$), whereas the subpopulation is allowed to have a different distribution of primary and secondary spin magnitudes ($\mu^{\chi_{1}}_{S}\neq\mu^{\chi_{2}}_{S}$, $\sigma^{\chi_{1}}_{S}\neq\sigma^{\chi_{2}}_{S}$), to allow for the 2G+1G scenario. We additionally fix $\mu^{\chi_{1}}_{S}=0.69$ and $\sigma_{S}^{\chi_{1}}=0.1$ to encourage the subpopulation to represent binaries with 2G primary components. Although this is a strong prior choice, we do not enforce the existence of this subpopulation, as $\xi$ is a free parameter which is allowed to be 0.
The mass distribution for both populations uses a pairing function formalism Fishbach and Holz (2020) with a Gaussian pairing function in mass ratio, $q=m_{2}/m_{1}$, truncated in the range $(0,1]$. Inspired by Tong et al. (2025b) and Farah et al. (2024), the hyperparameters governing both components are identical for the main population, except that we allow differing maximum masses, $m_{i}^{\max}$, for the primary and secondary components. For the subpopulation, all hyperparameters are allowed to differ between the primary and secondary masses as we expect 1G BHs to have a different mass distribution than 2G BHs. All component mass distributions are parametrized by the Broken Power Law + 2 Peaks primary mass distribution from The LIGO Scientific Collaboration et al. (2025). The resulting two-dimensional mass distribution is then
$p_{j}(m_{1},m_{2}|\lambda_{j})\propto\text{BP2P}(m_{1}|\lambda_{j}^{m_{1}})\text{BP2P}(m_{2}|\lambda_{j}^{m_{2}})\mathcal{N}(q|\mu^{q}_{j},\sigma^{q}_{j}).$ (A1)
For a description of hyperparameters governing the Broken Power Law + 2 Peaks model, we refer readers to Appendix B.3 of The LIGO Scientific Collaboration et al. (2025).
We require that the pairing function be identical for the main and subpopulations, which encodes our assumption that the 1G+1G mergers in our sample are subject to the same gravitational potential and therefore undergo the same process of mass segregation as 2G+1G mergers. However, this assumption might not be valid as star cluster potentials may be time-varying, or, alternatively, our observed 2G+1G mergers may come from a distinct cluster population from our observed 1G+1G mergers. We therefore relax this assumption in Appendix E, and find that our main conclusions are largely unchanged, although our mass ratio distributions differ noticeably with different pairing function assumptions.
Appendix B GW231123
Including GW231123 does not impact our redshift or spin results. However, it does impact our inference of the secondary mass distribution. When included, the event appears to be fully absorbed in the subpopulation, and causes the maximum secondary mass of the subpopulation to shift to large values that are inconsistent with those inferred when GW231123 is excluded. We therefore deem GW231123 an outlier in $m_{2}$ and exclude it from the analysis in the main text of this paper. One possible interpretation of GW231123’s inconsistency with the inferred secondary mass distribution of the subpopulation is that it is not a 2G+1G merger. However, we refrain from making definitive or quantitative statements on this event.
Appendix C Spin Distributions
C.1 Spin Orientations
Figure 5: Distribution of spin orientations for the main population and subpopulation. The subpopulation prefers isotropically-oriented spins, which is consistent with expectations for a cluster origin and further validates the interpretation of the subpopulation as representing 2G+1G systems. The main population has support for both an isotropic component and one peaked near aligned spins.
Fig. 5 shows the inferred spin orientation distributions for our main and subpopulation. Parameterized by $\cos{\theta_{i}}$, and assumed to be independent and identically distributed between the two black holes in each system, these distributions will be uniform if the spins of each BH are isotropically oriented with respect to their orbits’ angular momentum and peaked near unity if they tend to be aligned with the orbital angular momentum of the system. Isolated formation channels typically predict a preference for aligned-spin systems, while star clusters predict isotropic orientations. Thus, the fact that the supopulation exhibits a uniform spin orientation distribution is consistent with expectations from hierarchical mergers in dynamical environments.
The main population, however, has support for both isotropic and aligned-spin contributions, consistent with findings in The LIGO Scientific Collaboration et al. (2025). Aligned-spin systems nominally constitute $52^{+27}_{-25}\%$ of the main population, as indicated by the hyperposterior on the fraction of main-population systems contained within a Gaussian distribution centered at unity. While it is possible to interpret this as the percentage of isolated binary black hole mergers, the hyperparameter governing the fraction of systems in the Gaussian component is an over-estimate of the the fraction of aligned-spin systems, given that broad Gaussian distributions emulate uniform distributions.
In this work, we forego making definitive or quantitative statements from the spin orientation distribution, as GW observations have limited information on spin orientations and therefore population inference is subject to prior assumptions(Vitale et al., 2022a, 2025). However, the results in Fig. 5 encourage the interpretation of the subpopulation as a set of 2G+1G mergers in star clusters.
C.2 Spin Magnitudes
Figure 6: Distributions of spin magnitudes. The main population prefers low (but not necessarily zero) spins. The subpopulation’s primary spin magnitude distribution is fixed in shape but is allowed to vary in height, whereas its secondary spin magnitude distribution resembles its prior, so the data are not yet informative to the spin magnitudes of these systems.
The inferred spin magnitude distributions for our main and subpopulation are shown in Fig. 6. Remember that p($\chi_{1}$) for the subpopulation is a truncated Gaussian distribution with fixed mean and standard deviation, $\mu^{\chi_{1}}_{S}=0.69$, $\sigma_{S}^{\chi_{1}}=0.1$. All uncertainty in that distribution is therefore due to the hyperposterior on $\xi$, the mixture fraction between the main and subpopulation. The secondary spin magnitude distribution is fit to the data, and we find that it may be peaked between $0.08$ and $0.21$, indicating a preference for small but nonzero spin in the 1G components. However, the hyperposterior is broad and we are unable to draw definitive conclusions from current data. Furthermore, the spin magnitude hyperprior – shown in dashed lines in Fig. 6 – resembles the hyperposterior, so we caution against drawing conclusions from the secondary spin magnitude distribution of the subpopulation.
We assume that the two components of the main population’s systems are independent and identically distributed, and find them to be peaked strongly at $\mu_{M}^{\chi}=0.12^{+.04}_{-.05}$, again indicating a preference for small but nonzero spins, and consistent with The LIGO Scientific Collaboration et al. (2025).
Appendix D Mass ratio distributions
Figure 7: Mass ratio distributions, marginalized over hyperposterior uncertainty. The subpopulation’s mass ratio distribution is peaked at lower values than that of the main population, consistent with expectations for 2G+1G mergers. However, the value at which it peaks depends strongly on our choice of pairing function (see Fig. 9).
Our inferred mass ratio distributions are influenced both by our inferred primary and secondary component mass distributions, and by our inferred pairing function. The peak of our pairing function is inferred to be at $\mu^{q}=0.87^{+0.07}_{-.07}$. This results in the mass ratio distributions shown in Fig. 7. The subpopulation’s mass ratio distribution peaks at lower values than that of the main population, consistent with expectations for 2G+1G mergers.
However, the value at which it peaks depends strongly on our choice of pairing function. Assuming the same pairing function between the main and subpopulation – as we do in Fig. 7 and all other figures in the main text – results in a mass ratio distribution that peaks at $q=0.73$. Assuming a different pairing function between the main and subpopulations results in a mass ratio distribution that peaks at $\approx 0.5$, which is in line with conventional expectations from 2G+1G mergers (Rodriguez et al., 2019; Ye et al., 2025). However, the data are not yet able to constrain the pairing function of the subpopulation, so this result is likely spurious.
Appendix E Choice of pairing function for subpopulation
Figure 8: Inferred pairing functions for the subpopulations when these are allowed to differ. The main population’s pairing function resembles what is inferred when the pairing is enforced to be identical between the two subpopulations, while the subpopulaion’s pairing function resembles its prior (dashed lines).
Figure 9: Mass ratio (left panel) and two-dimensional mass (right panel) distributions for the main and subpopulations in the case where the pairing function is allowed to differ between the main and subpopulation. All distributions are averaged over population-level uncertainty and colors are the same as in all previous figures.
The assumed form of the subpopulation’s pairing function primarily affects the inferred mass and mass ratio distributions of the subpopulation. In the main text, we enforce that the main and subpopulation have the same pairing function in order to emulate a constant mass segregation process in star clusters. However, this might not be a valid assumption. For example, if 2G+1G mergers happen later in a cluster’s history than 1G+1G mergers, the cluster’s gravitational potential and the BHs that it contains may be different by the time the majority of 2G+1G mergers happen, the mass segregation process and therefore the pairing function may be different for 1G+1G versus 2G+1G mergers. This motivates us to explore a different pairing function for the subpopulation in this Appendix. We show the resulting inferred pairing functions in Fig. 8. While the main population’s pairing function (solid violet lines) is well-constrained and appears different from its prior (dashed violet lines), the subpopulation’s pairing function resembles its prior (dashed green lines). Furthermore, the main population’s pairing function in Fig. 8 strongly resembles the pairing function inferred from combining the main and subpopulation. It therefore may not be possible to meaningfully constrain the pairing of the 2G+1G population, further motivating our decision to enforce the same pairing between the main population and subpopulation in the main text.
In Fig. 10, we show the population-averaged mass ratio distribution and two-dimensional mass distribution for the case in which the pairing functions differ between the main population and subpopulation (compare to to Figs. 7 and 10). If the pairing functions differ, the subpopulation’s mass ratio distribution peaks at $0.5$, which is more in line with expectations from 2G+1G mergers, though we caution that these results are likely prior-driven.
Results relating to spin and redshift distributions are unchanged between the two choices of pairing functions considered here.
Appendix F Additional figures
Figure 10: Two-dimensional mass distributions of the main and subpopulation. Colors are the same as in all previous figures.
Figure 11: Marginal primary (top) and secondary (bottom) mass distributions. These produce similar conclusions to those presented in Section V.
Figure 12: 99th percentile of the primary (solid) and secondary (dashed) mass distributions for the main and subpopulations. These are calculated directly from Fig. 2. The main population’s merger rate declines above $46\,\mathrm{M}_{\odot}$ in both primary and secondary mass, whereas the subpopulation declines above $\approx 55\,\mathrm{M}_{\odot}$ in secondary mass and beyond $110\,\mathrm{M}_{\odot}$ in primary mass.
Fig 10 shows our inferred two-dimensional primary and secondary mass distributions of the main and subpopulations.
Fig. 11 shows the marginal component mass distributions, which differ from Fig. 2 because they include the effects of the pairing function, but the same qualitative features appear in both representations. The top panel of Fig. 11 can be directly compared to Fig. 3 of Tong et al. (2025a), and does exhibit qualitatively similar features. Specifically, the 2G+1G population peaks at $\approx 17\,\mathrm{M}_{\odot}$, may drop to negligible rates between $25\,\mathrm{M}_{\odot}$ and $30\,\mathrm{M}_{\odot}$, and does not require a secondary mode, instead preferring a shallow tail to high masses. The 1G+1G population is also consistent between Fig. 11 and Fig. 3 of Tong et al. (2025a), with a dip at $\approx 17\,\mathrm{M}_{\odot}$ and peak at $\approx 9\,\mathrm{M}_{\odot}$.
Fig. 12 displays the 99th percentiles of the component mass distributions in Fig. 2. It would seem to imply that the pair-instability mass feature begins at $\approx 46\,\mathrm{M}_{\odot}$. This is consistent with Wang et al. (2022), Pierra et al. (2024), and Antonini et al. (2025), but appears inconsistent with Wang et al. (2025), who find a population of low-spin objects extending to $\approx 70\,\mathrm{M}_{\odot}$. To directly compare to this finding, we calculate the fraction of systems in the spinning subpopulation in various mass bins and find that this fraction lower in the range $50\,\mathrm{M}_{\odot},70\,\mathrm{M}_{\odot}]$ than in the range $70\,\mathrm{M}_{\odot},90\,\mathrm{M}_{\odot}]$ to 91.56%. This supports the hypothesis put forth in Wang et al. (2025), but at low significance.

### v2

@@
-
+The steep redshift evolution of the hierarchical binary black hole merger rate may cause the $z$-$\chi_{\rm eff}$ correlation
 Abstract
@@
 I Introduction
-The stellar-mass binary black holes detectable by LIGO (LIGO Scientific Collaboration et al., 2015; Capote et al., 2025; Soni et al., 2025), Virgo (Acernese et al., 2015) and KAGRA (Akutsu et al., 2021) (LVK) are likely created from the collapse of massive stars.
-These stars may be born as a binary system in the galactic field that then evolves into a binary black hole system, or they may be born in a dense stellar environment such as a star cluster, in which the black hole stellar remnants dynamically assemble into tightly-bound binaries, and typically encounter other objects in the cluster before merger (see reviews by Mapelli, 2020; Mandel and Farmer, 2022, and references therein).
-While it is not yet known which of these potential formation scenarios, if any, contributes to the majority of binary black hole systems, one clear signature of the dynamical assembly hypothesis is the existence of hierarchical mergers (e.g.
+Star clusters are foundational to much of our understanding of galactic and stellar astrophysics, as they play a large role in galactic structure and dynamics (Harris and Racine, 1979), significantly contribute to the chemical evolution of their galaxies (Shara and Hurley, 2002), and can be used to trace galaxy assembly (Brodie and Strader, 2006).
+Crucial but uncertain features of the star cluster population, such as its mass function and its evolution with redshift (Forbes et al., 2018), can be probed by the stellar-mass binary black holes observed by the LIGO (LIGO Scientific Collaboration et al., 2015; Capote et al., 2025; Soni et al., 2025), Virgo (Acernese et al., 2015) and KAGRA (Akutsu et al., 2021) (LVK) gravitational-wave detectors (Fishbach and Fragione, 2023).
+However, this requires knowing which portion of the binary black hole population is produced by clusters.
+The clearest population-level signature of binary black hole formation in star clusters is the existence of hierarchical mergers (e.g.
 Gerosa and Fishbach, 2021).
-Hierarchical mergers include one or more “second-generation” remnants of previous, “first-generation” black hole mergers, and are informative for a number of reasons.
+These include one or more “second-generation” remnants of previous, “first-generation” black hole mergers.
+The number of mergers that include a “second-generation” black hole strongly informs the total rate of cluster-origin binary black holes.
+In this Letter, we investigate the population properties of the hierarchically-merged binary black holes, which can be used to inform the redshift and metallicity distributions of star clusters.
+Beyond their ability to identify which portion of the BBH population is sourced from clusters, hierarchical mergers are informative in their own right.
 Firstly, they are rarely produced in clusters with low stellar mass, as the recoil kicks that “second-generation” black holes receive from anisotropic emission of gravitational waves are typically larger than the clusters’ escape velocity.
@@
 The existence of a rapidly-spinning subpopulation with a steep redshift evolution provides a natural explanation of the previously-identified correlation between effective spin and redshift in the overall binary black hole population.
-Secondly, if assumptions about the typical star clusters that host dynamically-formed binary black holes are made, the number of mergers that include a “second-generation” black hole strongly informs the number of dynamically-formed “first-generation”+“first-generation” mergers in the population (Kimball et al., 2020), thereby constraining the total rate of dynamically assembled systems.
-Finally, hierarchical mergers have clear population predictions that allow them to be straightforwardly distinguished from the larger binary black hole population (Fishbach et al., 2017; Gerosa and Berti, 2017).
+Secondly, hierarchical mergers have clear population predictions that allow them to be straightforwardly distinguished from the larger binary black hole population (Fishbach et al., 2017; Gerosa and Berti, 2017).
 An ensemble of black holes that have resulted from previous mergers will have a spin magnitude distribution peaked near $\chi=0.69$ (Berti and Volonteri, 2008; Hofmann et al., 2016; Rodriguez et al., 2019; Borchers et al., 2025), regardless of how many previous mergers occurred (Fishbach et al., 2017).
@@
 Regardless, the fact that a highly-spinning subpopulation has a steeper redshift evolution than the rest of the binary black holes offers a potential explanation to the previously-discovered broadening of the effective spin distribution with redshift.
-In a companion paper (Vijaykumar et al., in prep.), we show that this subpopulation can also explain the effective spin distribution’s narrowing with mass ratio (Callister et al., 2021; The LIGO Scientific Collaboration et al., 2025).
+In a companion paper (Vijaykumar et al., 2026), we show that this subpopulation can also explain the effective spin distribution’s narrowing with mass ratio (Callister et al., 2021; The LIGO Scientific Collaboration et al., 2025).
 The low-mass end of our inferred 1G mass distribution differs between the 1G+1G and 2G+1G populations by more than would be expected from the simplest mass segregation models.
@@
 (2025), albeit with a different population model.
-Kimball et al.
+Pierra et al.
+(2024) find that $\approx 2$$5\%$ of the overall population belong to a $\chi\approx 0.7$ subpopulation, which is lower than that inferred in this work, potentially caused by differing inferred redshift distributions.
+Additionally, Kimball et al.
 (2021) find $\mathcal{R}_{\rm 2G+1G}/\mathcal{R}_{\rm 1G+1G}<0.12$ when analyzing GWTC-2, and our posterior on this value peaks at $0.23$, putting these two studies performed on different datasets in mild tension.
 (e.g.
-Kimball et al., 2021; Wang et al., 2022; Mould et al., 2022; Pierra et al., 2024; Antonini et al., 2025; Tong et al., 2025b, a; Berti et al., 2025; Vijaykumar et al., in prep.).
+Kimball et al., 2021; Wang et al., 2022; Mould et al., 2022; Pierra et al., 2024; Antonini et al., 2025; Tong et al., 2025b, a; Berti et al., 2025; Vijaykumar et al., 2026).
 The fact that similar fractions of 2G+1G systems are inferred – regardless of what compact binary parameter (i.e.
@@
 Note that the effective spin ($\chi_{\rm eff}$) distribution cannot be analytically calculated from our component spin model, so it is reconstructed from samples and is therefore shown as a histogram, marginalized over hyperparameter uncertainty.
+Additionally, the redshift distribution displayed in Fig.
+4 (bottom-right panel) is the (normalized) probability distribution over redshift, $p(z)$, – as opposed to the volumetric merger rate shown in Fig.
+1 – because $p(z)$ is directly obtained by marginalizing the two-dimensional probability distributions shown in the off-diagonal panels.
 All pairs of parameters shown in Fig.
@@
 (2022).
-In a companion paper (Vijaykumar et al., in prep.), we show that a subpopulation of hierarchical mergers also offers a natural explanation of the correlation first discovered in Callister et al.
+In a companion paper (Vijaykumar et al., 2026), we show that a subpopulation of hierarchical mergers also offers a natural explanation of the correlation first discovered in Callister et al.
 (2021), and then found by The LIGO Scientific Collaboration et al.
@@
 Using the prediction that second-generation BHs should have dimensionless spin magnitudes of $\chi\approx 0.69$, we identify a subpopulation of hierarchical mergers in the fourth gravitational-wave transient catalog.
-The merger rate and primary mass distribution of this subpopulation is consistent with those found previously with alternative methods (Tiwari and Fairhurst, 2021; Mould et al., 2022; Wang et al., 2022; Pierra et al., 2024; Antonini et al., 2025; Tiwari, 2025; Tong et al., 2025a; Vijaykumar et al., in prep.).
+The merger rate and primary mass distribution of this subpopulation is consistent with those found previously with alternative methods (Tiwari and Fairhurst, 2021; Mould et al., 2022; Wang et al., 2022; Pierra et al., 2024; Antonini et al., 2025; Tiwari, 2025; Tong et al., 2025a; Vijaykumar et al., 2026).
 Having isolated this hierarchically-formed subpopulation, we study its component mass, spin orientation, and redshift distributions.
@@
 If all observed correlations in the gravitational wave data can be attributed to a subpopulation of hierarchical mergers (see e.g.
-Vijaykumar et al., in prep.), a consistent picture of the binary black hole population is beginning to emerge.
+Vijaykumar et al., 2026), a consistent picture of the binary black hole population is beginning to emerge.
 This allows us to begin interpreting the rest of the binary black hole population, using the 2G+1G subpopulation as an anchor.
-Furthermore, the inferred population properties of hierarchical mergers allows us to understand the properties of star clusters from which they are sourced.
+The population properties of hierarchical mergers inferred in this work will enable constraints on the properties of star clusters from which they are sourced.
+For instance, interpreting the secondary mass distribution of the 2G+1G subpopulation as the birth mass distribution of BHs in star clusters can inform the metallicity content of these clusters, as well as the supernova mechanism.
+Additionally, the the 2G+1G subpopulation’s redshift distribution informs the formation times, metallicity content, and redshift-dependent mass function of star clusters.
 The authors gratefully acknowledge helpful conversations with Hui Tong, Daniel Holz, Utkarsh Mali, Mike Zevin, Sharan Banajiri, and Thomas Dent.
@@
 Although this is a strong prior choice, we do not enforce the existence of this subpopulation, as $\xi$ is a free parameter which is allowed to be 0.
-The mass distribution for both populations uses a pairing function formalism Fishbach and Holz (2020) with a Gaussian pairing function in mass ratio, $q=m_{2}/m_{1}$, truncated in the range $(0,1]$.
+The mass distribution for both populations uses a pairing function formalism (Fishbach and Holz, 2020) with a Gaussian pairing function in mass ratio, $q=m_{2}/m_{1}$, truncated in the range $(0,1]$.
 Inspired by Tong et al.
 (2025b) and Farah et al.
-(2024), the hyperparameters governing both components are identical for the main population, except that we allow differing maximum masses, $m_{i}^{\max}$, for the primary and secondary components.
-For the subpopulation, all hyperparameters are allowed to differ between the primary and secondary masses as we expect 1G BHs to have a different mass distribution than 2G BHs.
+(2024), the hyperparameters governing both components are identical for the main population, except that we allow differing maximum masses for the primary and secondary components.
+For the subpopulation, all hyperparameters are allowed to differ between the primary and secondary masses, as we expect 1G BHs to have a different mass distribution than 2G BHs.
 All component mass distributions are parametrized by the Broken Power Law + 2 Peaks primary mass distribution from The LIGO Scientific Collaboration et al.
 (2025).
-The resulting two-dimensional mass distribution is then
-$p_{j}(m_{1},m_{2}|\lambda_{j})\propto\text{BP2P}(m_{1}|\lambda_{j}^{m_{1}})\text{BP2P}(m_{2}|\lambda_{j}^{m_{2}})\mathcal{N}(q|\mu^{q}_{j},\sigma^{q}_{j}).$ (A1)
-For a description of hyperparameters governing the Broken Power Law + 2 Peaks model, we refer readers to Appendix B.3 of The LIGO Scientific Collaboration et al. (2025).
+For a given subpopulation, the resulting two-dimensional mass distribution is then
+$p_{j}(m_{1},m_{2}|\lambda_{j})\propto\text{BP2P}(m_{1}|\lambda_{j}^{m_{1}})\text{BP2P}(m_{2}|\lambda_{j}^{m_{2}})\mathcal{N}(q|\mu^{q}_{j},\sigma^{q}_{j}),$ (A1)
+where $j=[S,M]$ denotes the particular subpopulation.
+For a description of hyperparameters governing the Broken Power Law + 2 Peaks model, we refer readers to Appendix B.3 of The LIGO Scientific Collaboration et al.
+(2025).
 We require that the pairing function be identical for the main and subpopulations, which encodes our assumption that the 1G+1G mergers in our sample are subject to the same gravitational potential and therefore undergo the same process of mass segregation as 2G+1G mergers.
@@
 We therefore relax this assumption in Appendix E, and find that our main conclusions are largely unchanged, although our mass ratio distributions differ noticeably with different pairing function assumptions.
+Both subpopulations have redshift distributions of the form (Fishbach et al., 2018)
+$p_{j}(z|\kappa_{j})\propto\frac{{\rm d}V_{c}}{{\rm d}z}(1+z)^{\kappa_{j}-1}.$ (A2)
 Appendix B GW231123

### v3

@@
 Using the prediction that these hierarchically-merged black holes have dimensionless spin magnitudes of $\chi\approx 0.69$, we identify a subpopulation in the gravitational-wave data consistent with a hierarchical-merger origin in dense star clusters.
-This subpopulation’s primary mass distribution peaks at $16.7^{+3.1}_{-4.6}\,\mathrm{M}_{\odot}$, which is approximately twice as large as its secondary mass distribution’s mode ($8.0^{+29.7}_{-2.3}\,\mathrm{M}_{\odot}$), and its spin tilt distribution is consistent with isotropy.
+This subpopulation’s primary mass distribution peaks at $17.0^{+18.3}_{-4.4}\,\mathrm{M}_{\odot}$, which is approximately twice as large as its secondary mass distribution’s mode ($10.5^{+29.7}_{-4.7}\,\mathrm{M}_{\odot}$), and its spin tilt distribution is consistent with isotropy.
 Our inferred secondary mass distributions imply that isolated binary evolution may still be needed to explain the entirety of the $9\,\mathrm{M}_{\odot}$ peak.
-Surprisingly, we find that the rate of hierarchical mergers may evolve more steeply with redshift than the rest of the population ($98.5\%$ credibility): the fraction of all binary black holes that are hierarchically formed at $z=0.1$ is $0.05^{+0.05}_{-0.04}$, compared to $0.17^{+0.13}_{-0.12}$ at $z=1$.
+Surprisingly, we find that the rate of hierarchical mergers may evolve more steeply with redshift than the rest of the population ($98.0\%$ credibility): the fraction of all binary black holes that are hierarchically formed at $z=0.1$ is $0.03^{+0.05}_{-0.02}$, compared to $0.09^{+0.11}_{-0.07}$ at $z=1$.
 This provides an explanation for the previously-discovered broadening of the effective spin distribution with redshift.
@@
 We omit GW231123  (Abac et al., 2025b) as its large secondary mass (above $100\,M_{\odot}$) appears to be an outlier with respect to our inferred secondary mass distributions.
-Including this event does not affect our redshift or spin inferences.
+Including this event does not qualitatively affect our redshift or spin inferences, though it slightly reduces the statistical significance of the result shown in Fig.
+1, and it visibly changes the subpopulation’s secondary mass distribution.
 Appendix B further discusses the effects of GW231123 on our inferred mass distributions.
@@
 We find strong evidence for a subpopulation of $\chi_{1}=0.69$ systems.
-The fraction of systems in this subpopulation, $\xi$, is larger than $0.02$ at the $99.7\%$ level, representing a Bayes factor of $20.3$.
+The fraction of systems in this subpopulation, $\xi$, is larger than $0.01$ at the $99.9\%$ level, representing a Bayes factor of $49.3$.
 This subpopulation has spin orientation, mass ratio, and component mass distributions characteristic of hierarchical mergers in star clusters (see Appendices C.1, D, and Section V, respectively).
-Marginalized over the range $z=[0,2]$, our inferred value of $\xi=0.23^{+0.18}_{-0.17}$ is roughly consistent with previous studies that also identify a 2G+1G-like subpopulation that makes up roughly $12-20\%$ of the population [^1] Antonini et al.
+Marginalized over the range $z=[0,2]$, our inferred value of $\xi=0.13^{+0.16}_{-0.10}$ is roughly consistent with previous studies that also identify a 2G+1G-like subpopulation that makes up roughly $12-20\%$ of the population [^1] Antonini et al.
 (2025) estimate that $20\%$ of all mergers are 2G+1G, and Tong et al.
@@
 Additionally, Kimball et al.
-(2021) find $\mathcal{R}_{\rm 2G+1G}/\mathcal{R}_{\rm 1G+1G}<0.12$ when analyzing GWTC-2, and our posterior on this value peaks at $0.23$, putting these two studies performed on different datasets in mild tension.
+(2021) find $\mathcal{R}_{\rm 2G+1G}/\mathcal{R}_{\rm 1G+1G}<0.12$ when analyzing GWTC-2, which is in agreement with our posterior on this value, which peaks at $0.07$.
 (e.g.
@@
 (2025), is $3.2^{+0.94}_{-1.00}$, which lies between the values we infer for our main and subpopulations.
-The posterior on $\kappa_{\rm main}$ is less than that on $\kappa_{\rm sub}$ at $98.5\%$ credibility.
+The posterior on $\kappa_{\rm main}$ is less than that on $\kappa_{\rm sub}$ at $98.0\%$ credibility.
 There are therefore hints, but not yet definitive evidence, that hierarchical mergers represent a higher fraction of the binary black hole population at $z=1$ compared to $z=0$.
@@
 Assuming a cluster mass function that scales inversely with the square of the cluster mass, the metallicity-specific star formation history from Madau and Fragos (2017), and the catalog of Cluster Monte Carlo simulations from Kremer et al.
-(2020), we infer the fraction of all BBHs that originated in star clusters to be $0.32^{+0.35}_{-0.23}$ at $z=0.1$ and $1.23^{+0.93}_{-0.84}$ at $z=1$.
-This directly follows from our measured fraction of 2G+1G mergers in the total population at different redshifts, which is $0.05^{+0.05}_{-0.04}$ at $z=0.1$ and $0.17^{+0.13}_{-0.12}$ at $z=1$.
-Thus, if our assumptions about the redshift-dependent mass function of clusters are accurate, contributions from other formation scenarios may be necessary to explain the merger rate at low redshift, but clusters can explain the full merger rate at high redshift.
-Additionally, the fact that the fraction of cluster-origin BBHs may be larger than unity at $z=1$ means that denser clusters at high redshift may be required to explain our inferred fraction of 2G+1G mergers.
+(2020), we infer the fraction of all BBHs that originated in star clusters to be $0.18^{+0.30}_{-0.14}$ at $z=0.1$ and $0.74^{+0.89}_{-0.55}$ at $z=1$.
+This directly follows from our measured fraction of 2G+1G mergers in the total population at different redshifts, which is $0.03^{+0.05}_{-0.02}$ at $z=0.1$ and $0.09^{+0.11}_{-0.07}$ at $z=1$.
+Thus, if our assumptions about the redshift-dependent mass function of clusters are accurate, contributions from other formation scenarios are necessary to explain the merger rate at low redshift, but clusters can explain the full merger rate at high redshift.
+Additionally, the fact that the fraction of cluster-origin BBHs at $z=1$ has support above unity means that denser clusters at high redshift may be required to explain our inferred fraction of 2G+1G mergers.
 Theoretical uncertainties on cluster properties as a function of redshift remain uncertain, and future work will aim to constrain these with the subpopulation’s inferred redshift distribution (e.g.
@@
 after the pairing function has been factored out), but we provide marginalized component mass distributions in Appendix F.
-The hierarchical subpopulation’s primary mass distribution (representing 2G black holes) has a mode at $16.7^{+3.1}_{-4.6}\,\mathrm{M}_{\odot}$, which may be approximately twice as large as the mode of its secondary mass distribution ($8.0^{+29.7}_{-2.3}\,\mathrm{M}_{\odot}$).
+The hierarchical subpopulation’s primary mass distribution (representing 2G black holes) has a mode at $17.0^{+18.3}_{-4.4}\,\mathrm{M}_{\odot}$, which may be approximately twice as large as the mode of its secondary mass distribution ($10.5^{+29.7}_{-4.7}\,\mathrm{M}_{\odot}$).
 This is consistent with a 2G+1G interpretation of the subpopulation.
@@
 Taken together, these results further support the hypothesis proposed in Tong et al.
-(2025b) that the pair-instability mass gap is polluted primarily by 2G+1G mergers, and not by 2G+2G mergers.
+(2025b) that the pair-instability mass gap is polluted primarily by 2G+1G mergers rather than by 2G+2G mergers.
 However, it is possible that “second-generation”+“second-generation” mergers are present but not yet distinguishable from the rest of the population, and we cannot rule out their presence.
@@
 Ye and Fishbach (2024) shows that if the cluster mass function does not evolve with redshift, no significant difference is expected between the redshift evolution of 1G+1G versus hierarchical mergers (except at very early times before any “second-generation” black holes have formed).
-Our results that hierarchical mergers tend to occur at higher redshifts than “first-generation”+“first-generation” mergers, and that the fraction of hierarchical mergers may be as high as $0.17^{+0.13}_{-0.12}$ at $z=1$, may therefore imply that the cluster mass function does in fact evolve with redshift, and that hierarchical mergers originate in star clusters that are denser, more massive, and at higher redshift than the lower-redshift population that hosts the rest of the observed mergers (e.g.
+Our results that hierarchical mergers tend to occur at higher redshifts than “first-generation”+“first-generation” mergers, and that the fraction of hierarchical mergers may be as high as $0.09^{+0.11}_{-0.07}$ at $z=1$, may therefore imply that the cluster mass function does in fact evolve with redshift, and that hierarchical mergers originate in star clusters that are denser, more massive, and at higher redshift than the lower-redshift population that hosts the rest of the observed mergers (e.g.
 Mai et al., 2025).
@@
 Additionally, the the 2G+1G subpopulation’s redshift distribution informs the formation times, metallicity content, and redshift-dependent mass function of star clusters.
-The authors gratefully acknowledge helpful conversations with Hui Tong, Daniel Holz, Utkarsh Mali, Mike Zevin, Sharan Banajiri, and Thomas Dent.
+The authors gratefully acknowledge helpful conversations with Hui Tong, Daniel Holz, Utkarsh Mali, Mike Zevin, Sharan Banagiri, and Thomas Dent.
 We extend special thanks to Bart Ripperda for granting us access to the Bee workstation at CITA, where the computations for this work were performed.
@@
 Appendix B GW231123
-Including GW231123 does not impact our redshift or spin results.
-However, it does impact our inference of the secondary mass distribution.
-When included, the event appears to be fully absorbed in the subpopulation, and causes the maximum secondary mass of the subpopulation to shift to large values that are inconsistent with those inferred when GW231123 is excluded.
+Including GW231123 does not qualitatively impact our redshift or spin results, but it does lessen the statistical significance of the finding that the subpopulation has a steeper redshift evolution than the main population.
+Because GW231123 has high primary spins, it appears to be fully absorbed in the subpopulation.
+It’s redshift is also relatively low ($z\approx 0.4$), which causes the subpopulation’s inferred redshift distribution to flatten.
+Thus, we find $\kappa_{M}<\kappa_{S}$ to $88\%$ when GW231123 is included in the analysis, compared to 98% when it is excluded.
+Furthermore, including GW231123 in the analysis impacts our inferred secondary mass distribution.
+When included, the event causes the maximum secondary mass of the subpopulation to shift to large values that are inconsistent with those inferred when GW231123 is excluded.
 We therefore deem GW231123 an outlier in $m_{2}$ and exclude it from the analysis in the main text of this paper.
@@
 (2025).
-Aligned-spin systems nominally constitute $52^{+27}_{-25}\%$ of the main population, as indicated by the hyperposterior on the fraction of main-population systems contained within a Gaussian distribution centered at unity.
-While it is possible to interpret this as the percentage of isolated binary black hole mergers, the hyperparameter governing the fraction of systems in the Gaussian component is an over-estimate of the the fraction of aligned-spin systems, given that broad Gaussian distributions emulate uniform distributions.
+Aligned-spin systems nominally constitute $62^{+24}_{-25}\%$ of the main population, as indicated by the hyperposterior on the fraction of main-population systems contained within a Gaussian distribution centered at unity.
+While it is possible to interpret this as the percentage of isolated binary black hole mergers in the 1G+1G population, the hyperparameter governing the fraction of systems in the Gaussian component is typically an over-estimate of the fraction of aligned-spin systems, given that broad Gaussian distributions emulate uniform distributions.
 In this work, we forego making definitive or quantitative statements from the spin orientation distribution, as GW observations have limited information on spin orientations and therefore population inference is subject to prior assumptions (Vitale et al., 2022a, 2025).
@@
 All uncertainty in that distribution is therefore due to the hyperposterior on $\xi$, the mixture fraction between the main and subpopulation.
-The secondary spin magnitude distribution is fit to the data, and we find that it may be peaked between $0.08$ and $0.21$, indicating a preference for small but nonzero spin in the 1G components.
-However, the hyperposterior is broad and we are unable to draw definitive conclusions from current data.
-Furthermore, the spin magnitude hyperprior – shown in dashed lines in Fig.
-6 – resembles the hyperposterior, so we caution against drawing conclusions from the secondary spin magnitude distribution of the subpopulation.
-We assume that the two components of the main population’s systems are independent and identically distributed, and find them to be peaked strongly at $\mu_{M}^{\chi}=0.12^{+.04}_{-.05}$, again indicating a preference for small but nonzero spins, and consistent with The LIGO Scientific Collaboration et al.
+The secondary spin magnitude distribution is fit to the data, and we find that it may be peaked between $0.02$ and $0.23$, indicating a preference for small but not necessarily zero spin in the 1G components.
+However, the hyperposterior is broad and the spin magnitude hyperprior – shown in dashed lines in Fig.
+6 – resembles the hyperposterior for the subpopulation, so we further caution against drawing conclusions from the secondary spin magnitude distribution of the subpopulation.
+We assume that the two components of the main population’s systems are independent and identically distributed, and find them to be peaked strongly at $\mu_{M}^{\chi}=0.12^{+.06}_{-.06}$, again indicating a preference for small but nonzero spins, and consistent with The LIGO Scientific Collaboration et al.
 (2025).
@@
 Our inferred mass ratio distributions are influenced both by our inferred primary and secondary component mass distributions, and by our inferred pairing function.
-The peak of our pairing function is inferred to be at $\mu^{q}=0.87^{+0.07}_{-.07}$.
+The peak of our pairing function is inferred to be at $\mu^{q}=0.85^{+0.07}_{-.06}$.
 This results in the mass ratio distributions shown in Fig.
@@
 Assuming the same pairing function between the main and subpopulation – as we do in Fig.
-7 and all other figures in the main text – results in a mass ratio distribution that peaks at $q=0.73$.
+7 and all other figures in the main text – results in a mass ratio distribution for the subpopulation that peaks at $q=0.73$.
 Assuming a different pairing function between the main and subpopulations results in a mass ratio distribution that peaks at $\approx 0.5$, which is in line with conventional expectations from 2G+1G mergers (Rodriguez et al., 2019; Ye et al., 2025).
-However, the data are not yet able to constrain the pairing function of the subpopulation, so this result is likely spurious.
+However, the data are not yet able to constrain the pairing function of the subpopulation, so this result may be spurious.
 Appendix E Choice of pairing function for subpopulation
@@
 (2025), who find a population of low-spin objects extending to $\approx 70\,\mathrm{M}_{\odot}$.
-To directly compare to this finding, we calculate the fraction of systems in the spinning subpopulation in various mass bins and find that this fraction lower in the range $50\,\mathrm{M}_{\odot},70\,\mathrm{M}_{\odot}]$ than in the range $70\,\mathrm{M}_{\odot},90\,\mathrm{M}_{\odot}]$ to 91.56%.
+To directly compare to this finding, we calculate the fraction of systems in the spinning subpopulation in various mass bins and find that this fraction is lower in the range $m_{1}\in[50\,\mathrm{M}_{\odot},70\,\mathrm{M}_{\odot}]$ than in the range $m_{1}\in[70\,\mathrm{M}_{\odot},90\,\mathrm{M}_{\odot}]$ to 91.1%.
 This supports the hypothesis put forth in Wang et al.